OpenStudy (anonymous):

Divide and simplify. x^2-1/x divided by x-1/x+9

OpenStudy (anonymous):

Please show me how to do this step by step. Thank you.

OpenStudy (anonymous):

\[x^2-1/x/(x-1/x+9) = (x+1)(x-1)/x)*(x+9/x-1) = (x+1)(x+9)/x\]

OpenStudy (anonymous):

Got cut out \[(x+1)(x+9)/x\]